OpenStudy (anonymous):

There is a sales tax of $8 on an item that costs $82 before tax. A second item costs $123 before tax. What is the sales tax on the second item?

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (johnweldon1993):

Well we can just set up a ratio here \[\large \frac{costA}{taxA} = \frac{costB}{taxB}\] \[\large \frac{82}{8} = \frac{123}{x}\] and solve for 'x'

OpenStudy (anonymous):

i would of never thought to do that set up but is it 12?

OpenStudy (johnweldon1993):

So we would cross multiply here \[\large 82x = 123\times 8\] And that would show us that \[\large x = \frac{123 \times 8}{82}\] and yes you get 12 :) great job!

OpenStudy (anonymous):

one more question?

OpenStudy (johnweldon1993):

Sure :P

OpenStudy (anonymous):

On the blueprint of a house, 24 millimeters represents 5 meters. The length of the living room is 30 millimeters on the blueprint. What is the actual length of the living room? would I solve it just like the last one?

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (johnweldon1993):

sorry hun, helping someone else And yeah, set up a ratio again :)

OpenStudy (anonymous):

its okay and \[\frac{ 24 }{ 5 }=\frac{ 30 }{ x }\] that?

OpenStudy (johnweldon1993):

Perfect!

OpenStudy (anonymous):

150?

OpenStudy (johnweldon1993):

Forgot to divide by the 24 :P

OpenStudy (johnweldon1993):

Remember cross multiply \[\large \frac{24}{5}= \frac{30}{x}\] \[\large 24x = 150\] \[\large x = \frac{150}{24}\]

OpenStudy (anonymous):

6.25

OpenStudy (johnweldon1993):

There we go! :)

OpenStudy (anonymous):

Maya runs 7 miles in 50 minutes. At the same rate, how many miles would she run in 75 \[\frac{ 7 }{ 50 }=\frac{ 75 }{ x }\]minutes?

OpenStudy (anonymous):

@johnweldon1993 last one I swear

OpenStudy (johnweldon1993):

mmhmm sure ;P haha and nope not quite Remember you have to stay consistent with what you have on top and on bottom here you have \[\large \frac{miles}{minutes} = \frac{minutes}{miles}\] You need them to be the same on top and bottom so we have \[\large \frac{7}{50} = \frac{x}{75}\]

OpenStudy (anonymous):

oh I thought 75 had to be on top

OpenStudy (johnweldon1993):

Nope, if you did want 75 on top...you would need to 50 on top in the other one too, just have to stay consistent :)

OpenStudy (anonymous):

0.7?

OpenStudy (johnweldon1993):

not quite what I get...

OpenStudy (anonymous):

so what did I do wrong?

OpenStudy (johnweldon1993):

What did you do wrong? We have \[\large \frac{7}{50} = \frac{x}{75}\] cross multiply \[\large 50x = 75\times 7\] \[\large x = \frac{75 \times 7}{50}\]

OpenStudy (anonymous):

10.5

OpenStudy (johnweldon1993):

There ya go :)

OpenStudy (anonymous):

thank you so much(:

OpenStudy (johnweldon1993):

Anytime :)